OpenStudy (anonymous):

how do i verify cot (x- pi/2)=-tanx i dont know how to do this so pls be patient with me!

OpenStudy (jdoe0001):

hmmm \(\bf cot\left( x\cfrac{\pi }{2} \right) = -tan(x)?\)

OpenStudy (anonymous):

cot (x- pi/2)=-tanx

OpenStudy (anonymous):

sorry about that!

OpenStudy (jdoe0001):

k

OpenStudy (jdoe0001):

hmmm what's the \(\bf cos\left(\frac{\pi }{2} \right)\quad and \quad sin\left(\frac{\pi }{2} \right)\)

OpenStudy (anonymous):

what?

OpenStudy (jdoe0001):

ehehh, what are, check your Unit Circle, the \(\bf cos\left(\frac{\pi }{2} \right)\quad and \quad sin\left(\frac{\pi }{2} \right)?\)

OpenStudy (anonymous):

-1/2 and 0 ?

OpenStudy (jdoe0001):

-1/2 and 0? check your unit circle closer

OpenStudy (anonymous):

(0,1) ? i dont know i just started

OpenStudy (jdoe0001):

hmmm, well, \(\bf cos\left(\frac{\pi }{2} \right)=0\quad and \quad sin\left(\frac{\pi }{2} \right)=1\) so we know that much

OpenStudy (anonymous):

how is that relevant?

OpenStudy (jdoe0001):

\(\bf cot\left( x-\frac{\pi }{2} \right)\implies \cfrac{cos\left( x-\frac{\pi }{2} \right)}{sin\left( x-\frac{\pi }{2} \right)} \\ \quad \\ \cfrac{cos(x)cos\left(\frac{\pi }{2} \right)+sin(x)sin\left(\frac{\pi }{2} \right)}{sin(x)cos\left(\frac{\pi }{2} \right)-cos(x)sin\left(\frac{\pi }{2} \right)} \\ \quad \\ \cfrac{cos(x)\cdot {\color{brown}{ 0}}+sin(x)\cdot {\color{brown}{ 1}}}{sin(x)\cdot {\color{brown}{ 0}}-cos(x)\cdot {\color{brown}{ 1}}}\implies \cfrac{sin(x)}{-cos(x)}\implies ?\)

OpenStudy (anonymous):

wait wait wait how did you get that huge second fraction

OpenStudy (jdoe0001):

recall \(\bf \textit{Sum and Difference Identities} \\ \quad \\ sin({\color{brown}{ \alpha}} - {\color{blue}{ \beta}})=sin({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})- cos({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}}) \\ \quad \\ cos({\color{brown}{ \alpha}} - {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}}) + sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}})\)

OpenStudy (anonymous):

a=x and b=pi/2 ???

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so sin/-cos=-tan because tan=sin/cos right

OpenStudy (jdoe0001):

\(\bf \textit{Sum and Difference Identities} \\ \quad \\ sin({\color{brown}{ x}} - {\color{blue}{ \frac{\pi }{2}}})=sin({\color{brown}{ x}})cos({\color{blue}{ \frac{\pi }{2}}})- cos({\color{brown}{ x}})sin({\color{blue}{ \frac{\pi }{2}}}) \\ \quad \\ cos({\color{brown}{ x}} - {\color{blue}{ \frac{\pi }{2}}})= cos({\color{brown}{ x}})cos({\color{blue}{ \beta}}) + sin({\color{brown}{ x}})sin({\color{blue}{ \frac{\pi }{2}}})\)

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x)\cdot {\color{brown}{ 0}}+sin(x)\cdot {\color{brown}{ 1}}}{sin(x)\cdot {\color{brown}{ 0}}-cos(x)\cdot {\color{brown}{ 1}}}\implies \cfrac{sin(x)}{-cos(x)}\implies \cfrac{sin(x)}{-1\cdot cos(x)} \\ \quad \\ \cfrac{1}{-1}\cdot \cfrac{sin(x)}{cos(x)}\implies - \cfrac{sin(x)}{cos(x)}\implies -tan(x)\)

OpenStudy (anonymous):

thank you sooo much

OpenStudy (jdoe0001):

yw